# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 4.2 Sweeping over ancestors and advance split times
Following the procedure laid out in Sect. 2, we apply each perturbation $\{\omega_{m}\}_{m=1}^{M}$ to a collection of ancestor events $\{\mathbf{x}(t_{n}^{*})\}_{n=1}^{N}$ at a range of ASTs $\{t_{n}^{*}-A_{j}\}_{j=1}^{J}$. We set the number of ancestors, $N$ to whichever is smaller: the total number of cluster maxima (see Sect. 3) in the short DNS, or 32. The ASTs sampled are $\{A_{j}\}_{j=1}^{J=20}=\{2,4,\ldots,40\}$, with a two-day spacing chosen as roughly half the period of small fluctuations in $R(\mathbf{x}(t))$ (see Fig. 7).
@@ 5.1 Perturbed ensembles: case studies
Fig. 7 displays a small but representative sample of boosted ensembles at two target latitudes: (a) $y_{0}=\frac{38}{64}L$ and (b) $y_{0}=\frac{26}{64}L$, at the (southern, northern) edges of the (northern, southern) westerly jets, where meridional wind shear is (positive, negative). The ancestor’s intensity (black dashed curves) reach their respective peaks at times $t^{*}=(3760,2702)$. Note the differences in peak value and peak shape: the upper latitude has long-lasting, flat maxima and the lower latitude has brief, spiky maxima. In fact, by up-down symmetry, the two severity timeseries are statistically equivalent after reflection about $\frac{1}{2}$, hence the upper tail of one is the lower tail of the other.
@@
We show the perturbed intensities launched from three ASTs $A\in\{2,16,32\}$, colored (red, orange, blue) respectively. Following the split time, the ensemble members spread apart from the parent and from each other, achieving their own peak values (severities) that differ in both amplitude and timing from the ancestor, the discrepancies increasing with $A$. The red curves ($A=2$) replicate the ancestral peak very closely; the orange curves ($A=16$) peak at substantially higher or lower levels, and up to 2$\sim$ 2∼ 2 days earlier or later. Still, the orange peaks are clearly dynamically related to the ancestral peaks. This is no longer true for the blue curves ($A=32$), whose intensity peaks are widely scattered in time and systematically lower than the ancestors’ peaks.
@@
Besides these three selected ASTs, each descendant is charted in (a,b).i as a circle color-coded by AST, positioned vertically at its severity value and horizontally at its launch time. A corresponding star is plotted in (a,b).ii, positioned vertically at its severity value (on a zoomed-in scale) and horizontally at its peak timing (constrained by the “argmax drift” parameter $\delta t^{*}=5$ days, as explained in Sect. 2.1). We can see the transition of the $R^{*}$ ensemble from tightly clustered (for short AST) to roughly independent and climatologically distributed (for long AST), and in between there is a golden window of opportunity where severities can be both large and diverse. The optimal AST must balance these two objectives, a task akin to the exploitation-exploration tradeoff in Bayesian optimization and reinforcement learning (e.g., Yang et al.,, 2022). In this light, the two functionals defined in Eqs. (25) and (26) are candidate acquisition functions.
@@
Figure 7: Boosted ensembles of two selected events: (a) time $t^{*}=3760$ at latitude $y_{0}=\frac{38}{64}L$, and (b) time $t^{*}=2702$ at latitude $y_{0}=\frac{26}{64}L$. These are times when the intensity function $R(\mathbf{x}(t))$ from the long DNS (dashed black curves) achieved a peak value (horizontal black lines) above the threshold $\mu[(\frac{1}{2})^{5}]$ (horizontal gray lines). For each AST $A\in\{2,4,\ldots,40\}$, an ensemble of perturbed events (descendants) is launched at $t^{*}-A$, indexed by $m=1,\ldots,21$. For three selected ASTs $A=2,16,32$, the full timeseries $\{R_{m}(t)\}_{m=1}^{21}$ are shown in (a,b).i. The red-to-blue color scale indicates short-to-long ASTs. Each descendant achieves a different severity $R_{m}^{*}$ (peak intensity), indicated by circles in (a,b).i at $(-A,R_{m}^{*})$ for all values of $A$. The peaks also occur at different times $t_{m}^{*}$, indicated in (a,b).ii by stars at $(t_{m}^{*}-t^{*},R_{m}^{*})$, again for all $A$ and colored accordingly.
@@ 5.2 Relating severities to impulses: case studies
How should the response functions be parameterized? The simplest choice would be a linear model, often used in numerical weather prediction to optimize ensemble spread by perturbing in the most-effective directions, so-called singular vectors (Diaconescu and Laprise,, 2012). However, linear models are strictly valid only for infinitesimal perturbations, hence short lead times. Similar logic should apply when optimizing for severity instead of ensemble spread, and indeed we demonstrate below that the COAST tends to lie beyond the range where a linear model $\widehat{R}^{*}$ is valid. We therefore construct a quadratic model as well, and it turns out that this minor upgrade is sufficient. Future work with more complex dynamics and objectives may call for more elaborate response functions (orthogonal polynomials, Gaussian processes, and neural networks for example), but we adhere to quadratic models in this study as a proof of concept that is easy to construct and interpret, which we do in the following two figures.
@@
$+\theta_{3}\text{Re}\{{\omega}\}^{2}+\theta_{4}\text{Re}\{{\omega}\}\text{Im}\{{\omega}\}+\theta_{5}\text{Im}\{{\omega}\}^{2}$ $\theta_{3},\theta_{4},\theta_{5}\text{ fitted for quadratic model only.}$. (41)
@@
We use ordinary least squares regression on the $M=21$ sampled impulses $\{\omega_{m}\}_{m=1}^{M}$ and associated severities $\{R_{n,j,m}^{*}\}$, in addition to the non-perturbed ancestor ($\omega_{0}:=0$) with severity $R_{n,j,0}^{*}=R_{n}^{*}$. A different set of coefficients is calculated separately for each ancestor $n$ and AST $A_{j}$. The response functions for the same ancestor event as in Figs. 7b are visualized in Fig. 8, using (a) the two-dimensional response surfaces, (b) the true vs. fitted response values, (c) the overall slope, measured by the linear coefficient magnitudes, (d) the overall curvature, measured by the quadratic fit’s Hessian eigenvalues, and (e) the overall linear and quadratic skills, measured by via the coefficient of determination $R^{2}$. The response surface gradually transforms from a linear plane, to a curved hilltop, to a saddle, to a jagged landscape, as AST increases. Accordingly, the linear and then the quadratic model lose their skill. The quadratic model is slightly better than the linear model for this particular event, but substantially better when averaged across all events (see the forthcoming Fig. 9c.i).
@@
Figure 8: Row (a) represents impulses as in Fig. 6, but additionally shows the responses to them separately at six sampled ASTs, which increase from right to left (launch time $t^{*}-A$ increases left to right). At the shortest AST shown, $A=2$, the response function is clearly linear: the impulses above and left of center are marked by $+$, representing an increased severity, and those below and right of center are marked by $\bullet$, representing decreased severity, with marker sizes representing the magnitude of the change. Colored curves represent level sets of the fitted linear (cyan) and quadratic (orange) models, with (solid, dashed, dotted) contours to differentiate (positive, zero, negative) changes to $R^{*}$. As AST increases, the impulses causing higher and lower severities become more intertwined and less linearly separable, as the orange contours progressively bend and separate from the cyan contours. Row (b) displays the true vs. fit responses. Row (d) shows that the linear components $\theta_{1},\theta_{2}$ are estimated similarly (at least in magnitude) regardless of whether quadratic terms are also included. Row (d) shows that the quadratic model implies a local maximum (both eigenvalues nonpositive) for most of the range $A<26$, beyond which the landscape starts looking less like a hilltop and more like a saddle. Row (e) displays the coefficients of determination, $R^{2}$ (not to be confused with intensity $R$ or severity $R^{*}$, which fortunately we never need to square).
@@ 5.3 Conditional severity PDFs: case studies
Equipped with response functions, we can now construct conditional severity PDFs using Eq. (9), which are displayed in Fig. 9a. For the same ancestor as in Fig. 8 and the same six ASTs, we can see the relationship between actually sampled perturbed severities (red dots), fitted severity PDFs (colored curves, one color for each input scale $s$) evaluated at the bins with lower boundaries $\{\mu[(\frac{1}{2})^{k}]:k=5,\ldots,14\}$, and the climatological PDF (black curves). As AST increases from right to left, the severity PDFs morph from narrow spikes centered at the ancestor severity to long, extended lumps reaching far beyond the ancestor severity, and then recede below the threshold $\mu[(\frac{1}{2})^{5}]$. The PDF’s motion resembles a wave crashing onto a shallow beach, blanketing the sand, and then retreating, hitting the true COAST somewhere in the middle stages. But this general behavior is strongly modulated by the choice of scale $s$: red PDFs, representing the smallest scale $s=0.06$, are narrower and located closer to the ancestral severity (horizontal black line) for all ASTs, whereas blue PDFs, representing the largest scale $s=0.9$, spread out further as a result of giving more weight to bigger impulses. This underscores our claim that the input distribution, an arbitrary choice, merits sensitivity analysis, and so we carry it through the remaining steps.
@@ 5.4 AST selection criteria: case studies
Panels (b) display the criteria proposed in Sect. 2.3 that might help determine in which stage of “wave breaking” the severity PDF finds the COAST. The EI and TE criteria shown in panels b.(i,ii) both exhibit non-monotonic behavior by design, maximizing at COASTs denoted $A^{\pounds}[\text{EI}]$ and $A^{\pounds}[\text{TE}]$ (see Sect. 2.3). The AST dependence can be heuristically understood in light of the PDFs in Fig. 9a:
@@
  • At large AST past 25$\sim$ 25∼ 25 days, the PDFs have diminishing mass above $\mu$, let alone $R_{n}^{*}$, which zeros out most of the contributions to both EI and TE.
The COAST can change with the scale $s$: even though the overall shapes of TE and EI don’t change very much, the location of their maxima might. The TE, for example, peaks at $A=10$ days with $s=0.9$ but at $A=14$ days with $s=0.06$, which aligns with the findings from Finkel and O’Gorman, (2024) that stronger stochastic forcing (larger scale) shortens the COAST because ensembles spread faster. Fortunately, as Fig. 10 will corroborate, differences are small especially for $s\geq 0.24$.
@@
Fig. 9b.(iii,iv) display two versions of pattern correlation $\rho$, defined in Sect. 2.3 for an arbitrary field $F$: the “global correlation” $\rho[c]$ uses the whole two-dimensional upper-layer concentration field $F(x,y)=c_{1}(x,y)$, and the “local correlation” $\rho[c(\cdot,y_{0})]$ uses only the single-latitude transect $F(x)=c_{1}(x,y_{0})$ at the target latitude $y_{0}$. Both drop off steadily with AST, although local correlation fluctuates more due to averaging a smaller spatial region. The influence of scale enters at the ensemble-averaging step, where the $m$th member’s pattern correlation $\rho[F_{0},F_{m}]$ is weighted by $p(\omega_{m},s,W)$. Since smaller perturbations take longer to grow, smaller input scales lead to slower dropoff of $\rho$ with $A$—but only at short lead times, where errors are still tiny. Beyond $A\approx 6$ and 10 days for global and local correlations respectively, decorrelation proceeds at a similar rate for all scales. The nominal threshold $\epsilon^{2}=1-(\frac{3}{8})^{2}$ is marked in both.
@@
Figure 9: Output probability distributions of $R^{*}$ and ensemble dispersion indicators for the same case study as Fig. 8. (a) For six ASTs (same as Fig. 8), perturbed severities are displayed as dark red circles, and the unperturbed (ancestral) severity is marked with a horizontal black line. Colored curves show the output PDFs above $\mu=0.52$ as inferred from the quadratic regression, for a range of scales from 0.06 (red) to 0.9 (blue). Black curves represent the climatological tail PDF, as inferred from the long DNS, which the conditional PDFs should converge to at long $A$. (b.i) expected improvement and (b.ii) conditional entropy as a function of AST. Vertical bars mark the respective optimal ASTs, which may depend on the scale. (b.iii) local and (b.iv) global correlations as a measure of ensemble dispersion. The horizontal dashed lines are positioned at $1-(\frac{3}{8})^{2}$, corresponding to the rule of thumb from Finkel and O’Gorman, (2024). The vertical axes are stretched with a modified sigmoid to magnify numbers close to one and zero. (c.i) Coefficients of determination for linear (cyan) and quadratic (orange) regressions, averaged across ancestors. (c.(ii-v)) same quantities as in b.(i-iv) but averaged across ancestors, with only the largest and smallest scales shown. All error bars show truncated (upper, lower) means, i.e., the mean across (above-average, below-average) ancestors: $\mathbb{E}[X|X(>,<)\mathbb{E}X]$ for a random variable $X$.

### v2

@@ 4.2 Sweeping over ancestors and advance split times
Following the procedure laid out in Sect. 2, we apply each perturbation $\{\omega_{m}\}_{m=1}^{M}$ to a collection of ancestor events $\{\mathbf{x}(t_{n}^{*})\}_{n=1}^{N}$ at a range of ASTs $\{t_{n}^{*}-A_{j}\}_{j=1}^{J}$. We set the number of ancestors, $N$ to whichever is smaller: the total number of cluster maxima (see Sect. 3) in the short DNS, or 32. Considering all latitudes, the minimum $N$ was 14, the median was 22, and the maximum 32 was found at four latitudes including $y_{0}=\frac{26}{64}L$ which we consider in more depth. In the equal-cost comparisons to be shown later, we restrict $N$ to smaller values. The ASTs sampled are $\{A_{j}\}_{j=1}^{J=20}=\{2,4,\ldots,40\}$, with a two-day spacing chosen as roughly half the period of small fluctuations in $R(\mathbf{x}(t))$ (see Fig. 7).
@@ 5.1 Perturbed ensembles: case studies
Fig. 7 displays a small but representative sample of boosted ensembles at two target latitudes at the inner edges of the two eastward jets: (a) $y_{0}=\frac{38}{64}L$ and (b) $y_{0}=\frac{26}{64}$. The ancestors’ intensity (black dashed curves) reach their respective peaks at times $t^{*}=4152$ for (a) and $2702$ for (b). Note the differences in peak value and peak shape: the upper latitude has long-lasting, flat maxima and the lower latitude has brief, spiky maxima. The statistical properties at these two locations, both in Fig. 7 and in Fig. 3, are approximately equivalent after reflection about $\frac{1}{2}$ ($c\to 1-c$), meaning the upper tail of one resembles the lower tail of the other. This can be understood by the approximate north-south symmetry of the tracer’s dynamics imposed by Dirichlet boundary conditions.
@@
We show the perturbed intensities launched from three ASTs $A\in\{2,16,32\}$, colored (red, orange, blue) respectively. Following the split time, the ensemble members spread apart from the parent and from each other, achieving their own peak values (severities) that differ in both amplitude and timing from the ancestor, the discrepancies increasing with $A$. The red curves ($A=2$) replicate the ancestral peak very closely; the orange curves ($A=16$) peak at substantially higher or lower levels, and up to $\sim 2$ days earlier or later. Still, the orange peaks are clearly dynamically related to the ancestral peaks. This is no longer true for the blue curves ($A=32$), whose intensity peaks are widely scattered in time and systematically lower than the ancestors’ peaks.
@@
Besides these three selected ASTs, each descendant is charted in Fig. 7(a,b).i as a circle color-coded by AST, positioned vertically at its severity value and horizontally at its launch time. A corresponding star is plotted in Fig. 7(a,b).ii, positioned vertically at its severity value (on a zoomed-in scale) and horizontally at its peak timing (constrained by the “argmax drift” parameter $\delta t^{*}=5$ days, as explained in Sect. 2.1). We can see the transition of the $R^{*}$ ensemble from tightly clustered (for short AST) to roughly independent and climatologically distributed (for long AST), and in between there is a golden window of opportunity where severities can be both large and diverse. The optimal AST must balance these two objectives, a task akin to the exploitation-exploration tradeoff in Bayesian optimization and reinforcement learning (e.g., Yang et al., 2022). In this light, the two functionals defined in Eqs. (28) and (29) are candidate acquisition functions.
@@
Figure 7: Boosted ensembles of two selected events: (a) time $t^{*}=4152$ at latitude $y_{0}=\frac{38}{64}L$, and (b) time $t^{*}=2702$ at latitude $y_{0}=\frac{26}{64}L$. These are times when the intensity function $R(\mathbf{x}(t))$ from the long DNS (dashed black curves) achieved a peak value (horizontal dashed black lines) above the threshold $\mu[(\frac{1}{2})^{5}]$ (horizontal gray lines). For each AST $A\in\{2,4,\ldots,40\}$, an ensemble of perturbed events (descendants) is launched at $t^{*}-A$, indexed by $m=1,\ldots,21$. For three selected ASTs $A=2,16,32$, the full timeseries $\{R_{m}(t)\}_{m=1}^{21}$ are shown in (a,b).i. The red-to-blue color scale indicates short-to-long ASTs. Each descendant achieves a different severity $R_{m}^{*}$ (peak intensity), indicated by circles in (a,b).i at $(-A,R_{m}^{*})$ for all values of $A$. The peaks also occur at different times $t_{m}^{*}$, indicated in (a,b).ii by stars at $(t_{m}^{*}-t^{*},R_{m}^{*})$, again for all AA and colored accordingly.
@@ 5.2 Relating severities to impulses: case studies
How should the response functions be parameterized? The simplest choice would be a linear model, often used in numerical weather prediction to optimize ensemble spread by perturbing in the most-effective directions, so-called singular vectors (Diaconescu and Laprise, 2012). However, linear models are strictly valid only for infinitesimal perturbations, hence short lead times. Similar logic should apply when optimizing for severity instead of ensemble spread, and indeed we demonstrate below that the COAST tends to lie beyond the range where a linear model $\widehat{R}^{*}$ is valid. We therefore construct a quadratic model as well, and it turns out that this minor upgrade is sufficient. Future work with more complex dynamics and objectives may call for more elaborate response functions (orthogonal polynomials, Gaussian processes, and neural networks for example), but we adhere to quadratic models in this study as a proof of concept that is easy to construct and interpret, which we do in the following two figures.
@@
$\widehat{R}^{*}(\omega;\theta)$ $=\theta_{0}+\theta_{1}\text{Re}\{{\omega}\}+\theta_{2}\text{Im}\{{\omega}\}$ $\theta_{0},\theta_{1},\theta_{2}\text{ fitted for both linear and quadratic models}$ (43)
$+\theta_{3}\text{Re}\{{\omega}\}^{2}+\theta_{4}\text{Re}\{{\omega}\}\text{Im}\{{\omega}\}+\theta_{5}\text{Im}\{{\omega}\}^{2}$ $\theta_{3},\theta_{4},\theta_{5}\text{ fitted for quadratic model only.}$ (44)
@@
We use ordinary least squares regression on the $M=21$ sampled impulses $\{\omega_{m}\}_{m=1}^{M}$ and associated severities $\{R_{n,j,m}^{*}\}$, in addition to the non-perturbed ancestor ($\omega_{0}:=0$) with severity $R_{n,j,0}^{*}=R_{n}^{*}$. A different set of coefficients is calculated separately for each ancestor $n$ and AST $A_{j}$. The response functions for the same ancestor event as in Figs. 7b are visualized in Fig. 8, using (a) the two-dimensional response surfaces, (b) the true vs. fitted response values, (c) the overall slope, measured by the linear coefficient magnitudes, (d) the overall curvature, measured by the eigenvalues of the Hessian of the quadratic fit, and (e) the overall linear and quadratic skills, measured by the coefficient of determination $R^{2}$. The response surface gradually transforms from a linear plane, to a curved hilltop, to a saddle, to a jagged landscape, as AST increases. Accordingly, the linear and then the quadratic model lose their skill. The quadratic model is slightly better than the linear model for this particular event, but substantially better when averaged across all events (see the forthcoming Fig. 9c.i), and so we will use quadratic models only as $\widehat{R}^{*}$ in the tail estimators.
@@
Figure 8: The response of an extreme event to perturbations: magnitude, phase, and timing. The event is the same as in Fig. 7b. Row (a) represents impulses as in Fig. 6, but additionally shows the responses to them separately at six sampled ASTs (2, 10, 18, 24, 32, and 40 days, marked with vertical gray lines in c-e), which increase from right to left (launch time $t^{*}-A$ increases left to right). Horizontal and vertical scales are equal. At the shortest AST shown, $A=2$, the response function is clearly linear: the impulses above and left of center are marked by ++, representing an increased severity, and those below and right of center are marked by $\bullet$, representing decreased severity, with marker sizes representing the magnitude of the change. Colored curves represent level sets of the fitted linear (cyan) and quadratic (orange) models, with (solid, dashed, dotted) contours to differentiate (positive, zero, negative) changes to $R^{*}$. Row (b) displays the quality of these models by plotting true vs. fit responses (again, horizontal and vertical scales are equal). As AST increases, the impulses causing higher and lower severities become more intertwined and less linearly separable, as the orange contours progressively bend and separate from the cyan contours. Accordingly, the modeled linear response ceases to correlate with the true response. The modeled quadratic response has a slightly longer range of good quality, but also fails for AST $\gtrsim 26$ days. Row (c) shows that the linear components $\theta_{1},\theta_{2}$ are estimated similarly (at least in magnitude) regardless of whether quadratic terms are also included. Row (d) shows that the quadratic model implies a local maximum (both eigenvalues nonpositive) for most of the range $A<26$, beyond which the landscape starts looking less like a hilltop and more like a saddle. Row (e) displays the coefficients of determination, $R^{2}$ (not to be confused with intensity $R$ or severity $R^{*}$, which fortunately we never need to square).
@@ 5.3 Conditional severity PDFs: case studies
Equipped with response functions approximated by quadratic models, we can now construct conditional severity PDFs using Eq. (12), which are displayed in Fig. 9a. For the same ancestor as in Fig. 8 and the same six ASTs, we can see the relationship between actually sampled perturbed severities (red circles and lines), fitted severity PDFs (colored curves, one color for each input scale $s$) evaluated at the bins with lower boundaries $\{\mu[(\frac{1}{2})^{k}]:k=5,\ldots,14\}$, and the climatological PDF (black curves). As AST increases from right to left, the severity PDFs morph from narrow spikes centered at the ancestor severity to long, extended lumps reaching far beyond the ancestor severity, and then recede below the threshold $\mu[(\frac{1}{2})^{5}]$. The PDF’s motion resembles a wave crashing onto a shallow beach, blanketing the sand, and then retreating, hitting the true COAST somewhere in the middle stages. But this general behavior is strongly modulated by the choice of scale $s$: red PDFs, representing the smallest scale $s=0.06$, are narrower and located closer to the ancestral severity (horizontal black line) for all ASTs, whereas blue PDFs, representing the largest scale $s=0.9$, spread out further as a result of giving more weight to bigger impulses. This underscores our claim that the input distribution, an arbitrary choice, merits sensitivity analysis, and so we carry it through the remaining steps.
@@ 5.4 AST selection criteria: case studies
Figure 10 display the criteria proposed in Sect. 2.4 that might help determine in which stage of “wave breaking” the severity PDF finds the COAST. The EI and TE criteria shown in panels Fig. 10(a,b) both exhibit non-monotonic behavior by design, maximizing at COASTs denoted $A^{\text{EI}}$ and $A^{\text{TE}}$ (see Sect. 2.4). The AST dependence can be heuristically understood in light of the PDFs in Fig. 9:
@@
  • At large AST past $\sim 25$ days, the PDFs have diminishing mass above $\mu$, let alone above the ancestor severity $R_{n}^{*}$, which zeros out most of the contributions to both EI and TE.
The COAST can change with the scale $s$: even though the overall shapes of TE and EI don’t change very much, the location of their maxima might. Fortunately, we will find changes in scale for $s\gtrsim 0.24$ to have negligible impact.
@@
Fig. 10(c,d) display two versions of pattern correlation $\rho$, defined in Sect. 2.4 for an arbitrary field $F$: the “global correlation” $\rho[c]$ uses the whole two-dimensional upper-layer concentration field $F(x,y)=c_{1}(x,y)$, and the “local correlation” $\rho[c(\cdot,y_{0})]$ uses only the single-latitude transect $F(x)=c_{1}(x,y_{0})$ at the target latitude $y_{0}$. Both drop off steadily with AST, although local correlation fluctuates more due to averaging a smaller spatial region. The influence of perturbation scale $s$ enters at the ensemble-averaging step, where the $m$th member’s pattern correlation $\rho[F_{0},F_{m}]$ is weighted by $p(\omega_{m},s,W)$. Since smaller perturbations take longer to grow, smaller input scales lead to slower dropoff of $\rho$ with AA—but only at short lead times, where errors are still tiny. Beyond $A\approx 6$ and 10 days for global and local correlations respectively, decorrelation proceeds at a similar rate with respect to increasing AST for all scales. The nominal threshold $\rho=\sqrt{1-(\frac{3}{8})^{2}}$ is marked in both, and gives a similar AST for local and global correlations but generally longer than implied by EI or TE.
Figure 9: Severities and their conditional distributions for the same case study as Fig. 7b. For six ASTs (same as Fig. 8, decreasing from left to right), perturbed severities are displayed as dark red circles along a vertical line, and the unperturbed (ancestral) severity is marked with a horizontal black line. Colored curves and stars show the severity PDFs above $\mu=0.52$ as inferred from the quadratic regression, for a range of scales $s$ from 0.06 (red) to 0.9 (blue). Black curves with stars represent the climatological tail PDF, as inferred from the long DNS, which we will seek to estimate by combining conditional distributions over many ancestors (not just the single ancestor considered here).
@@
Figure 10: Ensemble dispersion indicators as a function of AST, again for the same case study as Fig. 7b: (a) expected improvement EI, (b) thresholded entropy TE, (c) local and (d) global correlations. Colors indicate input scales $s$, from small (red: $s=0.06$) to large (blue: $s=0.9$). In (a,b), vertical bars mark the respective optimal ASTs, which may depend on the scale. In (c,d), horizontal dashed lines are positioned at $1-(\frac{3}{8})^{2}$, corresponding to the rule of thumb from Finkel and O’Gorman (2024), and vertical axes are stretched with a modified sigmoid to magnify numbers close to one and zero.